# Multi-label Guided Soft Contrastive Learning for Efficient Earth Observation Pretraining

### v1

@@ IV-B Downstream tasks
We evaluate the pretrained backbones by linear probing and fine-tuning in 10 downstream tasks, including 6 land cover land use classification/segmentation datasets: BigEarthNet [42], BigEarthNet-SAR [54], EuroSAT [55], EuroSAT-SAR [22], fMoW-sentinel [18] and DFC2020 [56], and 4 multispectral datasets covering different applications from GEO-Bench [57]: m-so2sat, m-brick-kiln, m-cashew-plantation and m-SA-crop-type.
@@ V-A1 BigEarthNet and EuroSAT
TABLE I: Linear-probing / fine-tuning results on BigEarthNet-10% [42] and EuroSAT [55]. We report a comprehensive comparison with SOTA EO foundation models. MS/SAR/RGB represent data modalities. $\dagger$: SoftCon starts from DINO/DINOv2 which were trained on ImageNet/LVD-142M. *: SkySense employs a mixed architecture (totally 2B parameters) with ViT-L and Geo-context attention to encode MS images; 875K steps with batch size 240 roughly count to 1000 epochs. Left/right: linear/finetune. Best scores in bold.
@@
We first report SoftCon results on BigEarthNet-10% [42] and EuroSAT [55] which are most commonly evaluated by SOTA EO foundation models. As can be seen from Table I, our models outperform most of the existing works in all scenarios. Specifically on BigEarthNet, our ResNet50 improves over other works with the same backbone by a large margin, with 2.7%/1.6% increase on linear/fine-tuning compared to the current best model, and 6.2%/5.2% increase compared to SeCo [15]. Our linear probing result for the first time outperforms fully supervised training from scratch. Notably, our small ResNet50 also outperforms many existing large ViT models in both linear and fine-tuning, verifying the effectiveness of convolutional backbones.
@@ V-B Land cover segmentation
We report land cover semantic segmentation results on DFC2020 [56] as is shown in Table IV. We fine-tune DeepLabv3+ [65] for ResNet50 and UperNet [66] for ViT backbones. Promisingly, our ResNet50 outperforms MoCo-v2 [44] by 3%, and is also better than existing large ViTs. Our ViT-S outperforms CROMA [23] by 0.9%, which is further improved by our ViT-base with 0.3%.
@@
V-C Other domain-specific applications
TABLE V: Transfer results with frozen encoders on four Sentinel-2 tasks in GEO-Bench [57]. We report top-1 accuracy/mIoU for classification/segmentation, respectively.
@@ V-C Other domain-specific applications
Finally, we evaluate SoftCon on four other domain-specific applications with the corresponding Sentinel-2 dataset collections from GEO-Bench [57]. These include two classification datasets m-so2sat and m-brick-kiln, and two segmentation datasets m-cashew-plantation and m-SA-crop-type. We conduct frozen-encoder training following [25], with a linear classifier for classification tasks and UperNet decoder for segmentation tasks. The results are shown in Table V. We compare SoftCon with CROMA [23] and a recent work OFA-Net [25], and report the official fine-tuning (from timm [67]) results in [57] for reference. The table shows that SoftCon significantly outperforms both CROMA and OFA-Net on three tasks, while only slightly worse than OFA-Net on m-SA-crop-type. Notably, our results with frozen-encoder outperform the official results with full fine-tuning on the two segmentation datasets, and only slightly worse on m-so2sat.
@@ VI-1 SoftCon loss
We ablate ResNet50 results on the soft contrastive loss in Equation 4 which are shown in Table VI. In line with our explanation in Section III-B, SoftCon alone is worse than Contrast alone since the 10 Dynamic World classes are too coarse-grained to fully represent the real-world semantics. However, combining the two losses provides significant benefits. When we degrade the SoftCon to SupCon in Equation 2, the performance drops as expected, while still better than contrastive loss alone. These results verify the effectiveness of using existing free annotations to boost Earth observation pretraining. Additionally, we ablate the weighting parameter $\lambda$ in Equation 4 and suggest a best value of 0.1. Note all results in this ablation are without continual pretraining, which we will ablate next.
TABLE VI: Ablation study on the SoftCon loss and the weighting parameter.
@@ VI-2 Continual pretraining
We ablate the continual pretraining strategy in Table VII, which verifies the benefits of loading vision foundation models instead of pretraining from scratch. Also, we compare the performance of different vision models from ImageNet supervised weights to modern self-supervised weights. Interestingly, the stronger the vision model, the continual pretraining performance is also better. We report ImageNet top-1 linear scores of the corresponding models as a reference, which are well aligned with the cross-domain continual pretraining results.
@@
TABLE VII: Ablation study on the source models of continual pretraining.
@@ VI-3 Siamese masking
Finally, we ablate the Siamese masking strategy introduced mainly to boost the training efficiency. As can be seen from Table VIII, masking 50% of the patches can save almost half of the GPU memory without much performance drop. Interestingly, masking 20% of the patches can lead to even better results than no masking. This suggests that MAE-like random masking can be seen as an effective data augmentation strategy in contrastive learning.
TABLE VIII: Ablation study on the Siamese masking ratio.
@@ VII Conclusion
This work is the first to explore two important free resources beyond pure satellite imagery for efficient Earth observation pretraining: open land cover products and open vision foundation models. We build a large-scale multi-label dataset by matching an unsupervised pretraining dataset SSL4EO-S12 with Dynamic World land cover maps. To effectively utilize the multi-label annotations, we propose a novel soft contrastive learning method that allows soft matching between images with different label distributions. Meanwhile, we introduce strong vision models such as DINOv2 to a simple but flexible continual pretraining framework with Siamese masking. We successfully and efficiently train multispectral and SAR foundation models of both CNN and Transformer backbones that achieve SOTA performances in 9 out of 10 downstream tasks. Future work will explore more effective and flexible multimodal continual pretraining methods to build strong cross-domain cross-modal foundation models.
@@ A-A1 Data source
The Sentinel-2 images are from SSL4EO-S12 [9], a multi-modal multi-temporal dataset specifically designed for large-scale self-supervised learning. The dataset consists of 4-seasonal Sentinel-1/2 SAR-optical images from 251,079 non-overlapping locations in the world, covering a wide range of geographical and temporal diversities. The multi-label annotations are derived from Dynamic World [6], a near-real-time dataset that provides continuous pixel-level land cover monitoring in 9 semantic classes. The Dynamic World segmentation maps are aumatically generated by algorithms developped on high-quality training data. We integrate the noisy maps into rather accurate scene-level classification labels.
@@ A-B1 Label number distribution
fig. 6 shows the distribution of the number of labels within each image. About 17% images contain a single label, while 70% images contain 4 or more labels.
@@ A-B2 Class distribution
fig. 7 shows the distribution of the number of images for each class. The number of images is well balanced in 7 common classes, while flooded vegetation and snow/ice are less represented.
@@ A-B3 Season distribution
fig. 8 shows the distribution of the number of seasonal patches for each location. More than 40% locations have all 4 seasons, and more than 95% locations have at least 2 seasons.
@@ A-C Benchmark
We provide a preliminary benchmark on SSL4EO-S12-ML as a supervised multi-label classification task in table IX. We split the dataset into 80% training data and 20% testing data according to the non-overlapping locations, and report micro and macro mAP on the test split. We use 10% of the training data and the full testing data for the benchmark. The table shows that this multi-label classification task is rather easy to solve, with a supervised micro mAP reaching 98.2%. This is similar to the training metric, indicating the balanced dataset split and rather good label quality. In addition, we also test the dataset as a downstream task to evaluate pretrained models. As is shown in the table, consistent performances as other popular datasets in the main paper are observed: 1) pretrained models improve upon random initialization; 2) SoftCon outperforms existing models such as SSL4EO. Therefore, SSL4EO-S12-ML can be considered as a global multi-label classification benchmark dataset that complements existing datasets such as BigEarthNet [42] which covers only Europe.
TABLE IX: Benchmark results of SSL4EO-S12-ML dataset. We use 10% of training data.
@@ Appendix B Parameter Efficient Continual Pretraining
TABLE X: Linear probing mAP scores on BigEarthNet-10% with different pretraining strategies. We report pretraining with full SSL4EO-S12 data and 10% data. PEFT indicates directly transferring the DINOv2 model on BigEarthNet with PEFT techniques. *: SoftCon represents our SoftCon-style fully continual pretraining used in the main paper.
@@
table X presents linear probing results on BigEarthNet with different pretraining strategies. Firstly, as the table shows, straightforward PEFT from DINOv2 does provide benefits compared to random frozen encoder. However, such benefits are rather limited compared to in-domain pretraining. Secondly, compared to pretraining from scratch, SoftCon-style continual pretraining offers significant improvement, especially when the in-domain data is limited in size: continual pretraining on 10% data is comparable to pretraining full data from scratch. Thirdly, all PECP strategies provide reasonable benefits with only a small fraction of parameters trainable, outperforming direct PEFT. Among them, LoRA performs slighter better than others. Similar to full continual pretraining, PECP also remains the performance when the size of pretraining data is restricted. Lastly, all the simple PECP strategies are not close to full continual pretraining, restricting the practical usage. This preliminary study motivates us to use full continual pretraining for SoftCon in the main papaer. However, with more research towards advanced designs, we believe PECP bears potential for future work, as the flexible adapters can pave the way towards unified cross-domain foundation models.
@@ C-A Pretraining
We load ResNet50 weights from DINO [49] and ViT-S/14 and ViT-B/14 weights (without register) from DINOv2 [13], and conduct continual pretraining for 100 epochs. For ViTs, we randomly mask out 20% patches and send the remaining patches to the trainable encoder. All patches without masking are sent to the momentum encoder. Training is distributed in two nodes each with 4 A100 GPUs.
@@ C-B1 Downstream tasks
TABLE XI: Linear probing hyperparameters on BigEarthNet, EuroSAT and fMoW-S2.
@@
TABLE XII: Fine-tuning hyperparameters on BigEarthNet, EuroSAT and fMoW-S2.
@@
TABLE XIII: ResNet linear probing (left) and fine-tuning (right) hyperparameters on BigEarthNet and EuroSAT.
@@
TABLE XIV: DFC2020 fine-tuning hyperparameters.
DFC2020 (ResNet) DFC2020 (ViT)
Backbone ResNet50 ViT-B/14, ViT-S/14
Input size 224x224 224x224
Augmentation
ResizedCrop (0.5,2.0),
HorizontalFlip,
VerticalFlip
ResizedCrop (0.5,2.0),
HorizontalFlip,
VerticalFlip
Batch size 16 16
Learning rate 1e-3 5e-3
LR schedule poly poly
Optimizer SGD AdamW
Weight decay 5.00E-04 5.00E-02
Warm-up 0 1000
Iter 20K 40K
Head DeepLabv3+ UperNet
TABLE XV: Transfer learning hyperparameters on the four GEO-Bench datasets.

### v2

@@ IV-B Downstream tasks
We evaluate the pretrained backbones by linear probing and fine-tuning in 11 downstream tasks, including 6 land cover land use classification/segmentation datasets: BigEarthNet [53], BigEarthNet-SAR [65], EuroSAT [66], EuroSAT-SAR [28], fMoW-sentinel [24] and DFC2020 [67], one change detection dataset OSCD [68], and 4 multispectral datasets covering different applications from GEO-Bench [69]: m-so2sat, m-brick-kiln, m-cashew-plantation and m-SA-crop-type.
@@
  • OSCD: a binary change detection dataset consisting of 24 pairs of multispectral images distributed worldwide. We use the official split: 14 training and 10 test pairs.
@@ V-A1 BigEarthNet and EuroSAT
TABLE I: Linear-probing / fine-tuning results on BigEarthNet-10% [53] and EuroSAT [66]. We report a comprehensive comparison with SOTA EO foundation models. MS/SAR/RGB represent data modalities. $\dagger$: SoftCon starts from DINO/DINOv2 which were trained on ImageNet/LVD-142M. # indicates ”the number of”. *: SkySense employs a mixed architecture (in total 2B parameters) with ViT-L and Geo-context attention to encode MS images; 875K steps with batch size 240 roughly count to 1000 epochs. Left/right: linear/finetune. Best scores in bold.
@@
We first report SoftCon results on BigEarthNet-10% [53] and EuroSAT [66] which are most commonly evaluated by SOTA EO foundation models. As can be seen from Table I, our models outperform most of the existing works in all scenarios. Specifically on BigEarthNet, our ResNet50 improves over other works with the same backbone by a large margin, with 2.7%/1.6% increase on linear/fine-tuning compared to the current best model, and 6.2%/5.2% increase compared to SeCo [21]. Our linear probing result for the first time outperforms fully supervised training from scratch. Notably, our small ResNet50 also outperforms many existing large ViT models in both linear and fine-tuning, verifying the effectiveness of ConvNet backbones.
@@ V-B Land cover segmentation
We report land cover semantic segmentation results on DFC2020 [67] as is shown in Table IV. We fine-tune DeepLabv3+ [77] for ResNet50 and UperNet [78] for ViT backbones. Promisingly, our ResNet50 outperforms MoCo-v2 [55] by 3%, and is also better than existing large ViTs. Our ViT-S outperforms CROMA [29] by 0.9%, which is further improved by our ViT-base with 0.3%. For visual comparison, we plot some example segmentation maps in Figure 6. As it shows, SoftCon captures more accurate and more fine-grained semantic information compared to SSL4EO-S12 [10].
@@
Figure 6: Example segmentation maps on the DFC2020 [67] dataset.
V-C Change detection
We report binary change detection results on OSCD [68] as is shown in Table Table V. We freeze the backbone and fine-tune a simple U-Net [79] for segmentation. The differences in feature maps between two timestamps are input to the network. SoftCon outperforms SeCo and SSL4EO in both recall and F1-score. The low precision score is due to the significant class unbalance, i.e.: predicting all pixels as unchanged would result in a good precision score. The combined F1-score highlights the superior performance of SoftCon.
TABLE V: Results with frozen encoders on OSCD [68].
Backbone Precision Recall F1
Rand. Init. RN50 72.3 13.8 23.1
SeCo [21] RN50 74.9 17.5 28.3
SSL4EO [10] RN50 70.2 23.4 35.1
SoftCon (ours) RN50 66.6 29.2 40.6
V-D Other domain-specific applications
TABLE VI: Transfer results with frozen encoders on four Sentinel-2 tasks in GEO-Bench [69]. We report top-1 accuracy/mIoU for classification/segmentation, respectively.
@@ V-D Other domain-specific applications
Finally, we evaluate SoftCon on four other domain-specific applications with the corresponding Sentinel-2 dataset collections from GEO-Bench [69]. These include two classification datasets m-so2sat and m-brick-kiln, and two segmentation datasets m-cashew-plantation and m-SA-crop-type. We conduct frozen-encoder training following [31], with a linear classifier for classification tasks and UperNet decoder for segmentation tasks. The results are shown in Table VI. We compare SoftCon with CROMA [29] and a recent work OFA-Net [31], and report the official fine-tuning (from timm [80]) results in [69] for reference. The table shows that SoftCon significantly outperforms both CROMA and OFA-Net on three tasks, while only slightly worse than OFA-Net on m-SA-crop-type. Notably, our results with frozen-encoder outperform the official results with full fine-tuning on the two segmentation datasets, and only slightly worse on m-so2sat.
@@ VI-1 SoftCon loss
We ablate ResNet50 results on the soft contrastive loss in Equation 4 which are shown in Table VII. In line with our explanation in Section III-B, SoftCon alone is worse than Contrast alone since the 10 Dynamic World classes are too coarse-grained to fully represent the real-world semantics. However, combining the two losses provides significant benefits. When we degrade the SoftCon to SupCon in Equation 2, the performance drops as expected, while still better than contrastive loss alone. These results verify the effectiveness of using existing free annotations to boost Earth observation pretraining. Additionally, we ablate the weighting parameter $\lambda$ in Equation 4 and suggest a best value of 0.1. Note all results in this ablation are without continual pretraining, which we will ablate next.
TABLE VII: Ablation study on the SoftCon loss and the weighting parameter.
@@ VI-2 Continual pretraining
We ablate the continual pretraining strategy in Table VIII, which verifies the benefits of loading vision foundation models instead of pretraining from scratch. Also, we compare the performance of different vision models from ImageNet supervised weights to modern self-supervised weights. Interestingly, the stronger the vision model, the continual pretraining performance is also better. We report ImageNet top-1 linear scores of the corresponding models as a reference, which are well aligned with the cross-domain continual pretraining results.
@@
TABLE VIII: Ablation study on the source models of continual pretraining. cont. indicates continual pretraining.
@@ VI-3 Siamese masking
Finally, we ablate the Siamese masking strategy introduced mainly to boost training efficiency. As can be seen from Table IX, masking 50% of the patches can save almost half of the GPU memory without much performance drop. Interestingly, masking 20% of the patches can lead to even better results than no masking. This suggests that MAE-like random masking can be seen as an effective data augmentation strategy in contrastive learning. From another perspective, this design implicitly introduces the idea of masked image modeling that the model needs to know what information the masked patches have according to the unmasked branch, after which it knows the encoded features are similar. This could potentially also be one reason why the ideal masking ratio is much smaller compared to MAE, as the implicit masked image modeling plus the contrastive learning is a more challenging optimization task.
In summary, Table X provides an overview of the performance gains of each of our proposed components. It is worth noting that when introducing the multi-label annotation, the supervised contrastive learning method is implicitly introduced, which can be improved by our proposed soft contrastive learning as shown by the ablation in Table VII. Furthermore, when using the continual pretraining method, the vision dataset (e.g. ImageNet) is implicitly used for the vision model weights.
TABLE IX: Ablation study on the Siamese masking ratio.
@@
TABLE X: Performance gains of different components under linear probing on BigEarthNet-10%. The explicitly introduced components are highlighted in bold.
Method Backbone Dataset Performance gain
Contrast RN50 SSL4EO-S12 -
[…]
+ SoftCon RN50 - +0.6
+ cont. RN50 (+ ImageNet) +1.2
+ mask ViT-S - +0.5
@@ VII Conclusion
This work revisits two important free resources beyond pure satellite imagery for efficient Earth observation pretraining: open land cover products and open vision foundation models. We build a large-scale multi-label dataset by matching an unsupervised pretraining dataset SSL4EO-S12 with Dynamic World land cover maps. To effectively utilize the multi-label annotations, we propose a novel soft contrastive learning method that allows soft matching between images with different label distributions. Meanwhile, we introduce strong vision models such as DINO and DINOv2 to a simple but flexible continual pretraining framework with Siamese masking. We efficiently train multispectral and SAR foundation models of both CNN and Transformer backbones that achieve SOTA performances in 10 out of 11 downstream tasks.
There are two main limitations of this work. First, our models target mainly medium-resolution data, while high-resolution images with more fine-grained semantic information remain to be explored. Second, retraining vision foundation models without any constraints inevitably makes the model forget knowledge from the source vision domain. Future work will explore more effective and flexible multimodal continual pretraining methods to build strong cross-domain cross-modal foundation models.
@@ A-A1 Data source
The Sentinel-2 images are from SSL4EO-S12 [10], a multi-modal multi-temporal dataset specifically designed for large-scale self-supervised learning. The dataset consists of 4-seasonal Sentinel-1/2 SAR-optical images from 251,079 non-overlapping locations in the world, covering a wide range of geographical and temporal diversities. The multi-label annotations are derived from Dynamic World [6], a near-real-time dataset that provides continuous pixel-level land cover monitoring in 9 semantic classes. The Dynamic World segmentation maps are automatically generated by algorithms developed on high-quality training data. We integrate the noisy maps into rather accurate scene-level classification labels.
@@ A-B1 Label number distribution
Figure 7 shows the distribution of the number of labels within each image. About 17% images contain a single label, while 70% images contain 4 or more labels.
@@ A-B2 Class distribution
Figure 8 shows the distribution of the number of images for each class. The number of images is well balanced in 7 common classes, while flooded vegetation and snow/ice are less represented.
@@ A-B3 Season distribution
Figure 9 shows the distribution of the number of seasonal patches for each location. More than 40% locations have all 4 seasons, and more than 95% locations have at least 2 seasons.
@@ A-C Benchmark
We provide a preliminary benchmark on SSL4EO-S12-ML as a supervised multi-label classification task in Table XI. We split the dataset into 80% training data and 20% testing data according to the non-overlapping locations, and report micro and macro mAP on the test split. We use 10% of the training data and the full testing data for the benchmark. The table shows that this multi-label classification task is rather easy to solve, with a supervised micro mAP reaching 98.2%. This is similar to the training metric, indicating the balanced dataset split and rather good label quality. In addition, we also test the dataset as a downstream task to evaluate pretrained models. As is shown in the table, consistent performances as other popular datasets in the main paper are observed: 1) pretrained models improve upon random initialization; 2) SoftCon outperforms existing models such as SSL4EO. Therefore, SSL4EO-S12-ML can be considered as a global multi-label classification benchmark dataset that complements existing datasets such as BigEarthNet [53] which covers only Europe.
TABLE XI: Benchmark results of SSL4EO-S12-ML dataset. We use 10% of training data.
@@ A-D Data examples
Figure 10 provides some example images and corresponding multi-labels in the SSL4EO-S12-ML dataset covering different landscapes.
@@ Appendix B Parameter Efficient Continual Pretraining
TABLE XII: Linear probing mAP scores on BigEarthNet-10% with different pretraining strategies. We report pretraining with full SSL4EO-S12 data and 10% data. PEFT indicates directly transferring the DINOv2 model on BigEarthNet with PEFT techniques. *: SoftCon represents our SoftCon-style fully continual pretraining used in the main paper.
@@
Table XII presents linear probing results on BigEarthNet with different pretraining strategies. Firstly, as the table shows, straightforward PEFT from DINOv2 does provide benefits compared to random frozen encoder. However, such benefits are rather limited compared to in-domain pretraining. Secondly, compared to pretraining from scratch, SoftCon-style continual pretraining offers significant improvement, especially when the in-domain data is limited in size: continual pretraining on 10% data is comparable to pretraining full data from scratch. Thirdly, all PECP strategies provide reasonable benefits with only a small fraction of parameters trainable, outperforming direct PEFT. Among them, LoRA performs slighter better than others. Similar to full continual pretraining, PECP also remains the performance when the size of pretraining data is restricted. Lastly, all the simple PECP strategies are not close to full continual pretraining, restricting the practical usage. This preliminary study motivates us to use full continual pretraining for SoftCon in the main papaer. However, with more research towards advanced designs, we believe PECP bears potential for future work, as the flexible adapters can pave the way towards unified cross-domain foundation models.
@@ C-A Pretraining
We load ResNet50 weights from DINO [60] and ViT-S/14 and ViT-B/14 weights (without register) from DINOv2 [16], and conduct continual pretraining for 100 epochs. For ViTs, we randomly mask out 20% patches and send the remaining patches to the trainable encoder. All patches without masking are sent to the momentum encoder. Training is distributed in two nodes each with 4 NVIDIA A100 GPUs. Detailed compute and training time for different backbones are shown in Table XIII.
TABLE XIII: Compute and pretraining time.
Modality Backbone GPUs Training time
MS RN50 4xA100 21h
ViT-S/14 4xA100 25h
ViT-B/14 8xA100 15h
[…]
ViT-S/14 4xA100 8h
ViT-B/14 8xA100 7h
@@ C-B1 Downstream tasks
TABLE XIV: DFC2020 fine-tuning hyperparameters.
DFC2020 (ResNet) DFC2020 (ViT)
Backbone ResNet50 ViT-B/14, ViT-S/14
Input size 224x224 224x224
Augmentation
ResizedCrop (0.5,2.0),
HorizontalFlip,
VerticalFlip
ResizedCrop (0.5,2.0),
HorizontalFlip,
VerticalFlip
Batch size 16 16
Learning rate 1e-3 5e-3
LR schedule poly poly
Optimizer SGD AdamW
Weight decay 5.00E-04 5.00E-02
Warm-up 0 1000
Iter 20K 40K
Head DeepLabv3+ UperNet
TABLE XV: Linear probing hyperparameters on BigEarthNet, EuroSAT and fMoW-S2.
@@
TABLE XVI: Fine-tuning hyperparameters on BigEarthNet, EuroSAT and fMoW-S2.
@@
TABLE XVII: ResNet linear probing (left) and fine-tuning (right) hyperparameters on BigEarthNet and EuroSAT.
@@
TABLE XVIII: Transfer learning hyperparameters on the four GEO-Bench datasets.
@@
[Uncaptioned image] Yi Wang (S’21) received his B.E. degree in remote sensing science and technology from Wuhan University, Wuhan, China, in 2018 and his M.Sc. degree in geomatics engineering from University of Stuttgart, Stuttgart, Germany, in 2021. He is pursuing his Ph.D. degree at the Technical University of Munich (TUM), Munich, Germany. From 2021 to 2024, he was a research associate at the Remote Sensing Technology Institute, German Aerospace Center (DLR). In 2020, he spent three months at the perception system group, Sony Corporation, Stuttgart, Germany. His research interests include self-supervised learning, weakly-supervised learning, and multimodal representations.
[Uncaptioned image] Conrad M Albrecht (M’17) is a researcher at the Earth Observation Center of the German Aerospace Center. Since April 2021, he is PI of the HelmholtzAI Young Investigator Group (YIG) ”Large-Scale Data Mining in Earth Observation” (DM4EO). In July 2023, he was appointed a visiting associate professor with the Institute of Nasca at Yamagata University, Japan contributing to research in machine learning for the UNESCO World Heritage of the Nasca culture in Peru. For over 6 years Conrad was a research scientist in the Physical Sciences department at the IBM T.J. Watson Research Center in Yorktown, NY, USA. While at the Institute for Theoretical Physics, he graduated in physics (International Max-Planck Research School for Quantum Dynamics in Physics, Chemistry and Biology) with an extra certification in computer science (Cluster- & Detector Management team at CERN, Switzerland). He received a corresponding Ph.D. degree from Heidelberg University, Germany in 2014 working on distributed computing to study physics at low temperatures. Conrad’s research agenda interconnect physical models and numerical analysis, employing Big Data technologies and machine learning through open-science research, https://conrad-m-albrecht.github.io. Conrad co-organized workshops at the IEEE BigData conference, IGARSS conferences, and the AAAS annual meeting. Conrad is home in Europe and the United States. Some of his transatlantic initiatives include: In 2023, he was invited by the Alexander-von-Humboldt Foundation to present at the German-American Frontiers of Engineering Symposium. Beginning in 2024, initiated by Helmholtz Imaging and Data Science, he is in close touch with the German Department at Princeton University for cross-atlantic undergraduate internships. For fall 2024, Conrad was invited Adjunct Professor of Earth and Environmental Engineering by Columbia University, NY, USA.
[Uncaptioned image] Xiao Xiang Zhu (S’10–M’12–SM’14–F’21) received the Master (M.Sc.) degree, her doctor of engineering (Dr.-Ing.) degree and her “Habilitation” in the field of signal processing from Technical University of Munich (TUM), Munich, Germany, in 2008, 2011 and 2013, respectively. She is the Chair Professor for Data Science in Earth Observation at Technical University of Munich (TUM) and was the founding Head of the Department “EO Data Science” at the Remote Sensing Technology Institute, German Aerospace Center (DLR). Since May 2020, she is the PI and director of the international future AI lab ”AI4EO – Artificial Intelligence for Earth Observation: Reasoning, Uncertainties, Ethics and Beyond”, Munich, Germany. Since October 2020, she also serves as a Director of the Munich Data Science Institute (MDSI), TUM. From 2019 to 2022, Zhu has been a co-coordinator of the Munich Data Science Research School (www.mu-ds.de) and the head of the Helmholtz Artificial Intelligence – Research Field “Aeronautics, Space and Transport”. Prof. Zhu was a guest scientist or visiting professor at the Italian National Research Council (CNR-IREA), Naples, Italy, Fudan University, Shanghai, China, the University of Tokyo, Tokyo, Japan and University of California, Los Angeles, United States in 2009, 2014, 2015 and 2016, respectively. She is currently a visiting AI professor at ESA’s Phi-lab, Frascati, Italy. Her main research interests are remote sensing and Earth observation, signal processing, machine learning and data science, with their applications in tackling societal grand challenges, e.g. Global Urbanization, UN’s SDGs and Climate Change. Dr. Zhu has been a member of young academy (Junge Akademie/Junges Kolleg) at the Berlin-Brandenburg Academy of Sciences and Humanities and the German National Academy of Sciences Leopoldina and the Bavarian Academy of Sciences and Humanities. She is a Fellow of the Academia Europaea (the Academy of Europe). She serves in the scientific advisory board in several research organizations, among others the German Research Center for Geosciences (GFZ, 2020-2023) and Potsdam Institute for Climate Impact Research (PIK). She is an associate Editor of IEEE Transactions on Geoscience and Remote Sensing, Pattern Recognition and served as the area editor responsible for special issues of IEEE Signal Processing Magazine (2021-2023). She is a Fellow of IEEE, AAIA, and ELLIS.